What is the probability of choosing two hearts and then choosing two cards less then five, when (Ace=1)
Answer by stanbon(75887)

What is the probability of choosing two hearts and then choosing two cards less then five, when (Ace=1)
---
There are 16 cards less than 5 and 4 of them are hearts.
There are 16C2 = 120 pairs of cards less than 5.
There are 13C2 = 78 pairs of hearts
There are 4C2= 6 pairs of cards that are <5 and hearts
--------------------------
P(2H and 2(<5)) = P(2(<5))|2H)*P(2H)
= [6/78][78/52C2] = 6/52C2
= 6/1326
=0.0045
--------------
Cheers,
Stan H.